OpenStudy (anonymous):

Find the derivative of: y = cos^2(x) + cos(x^2)

OpenStudy (anonymous):

y = (cos(x))^2 + cos(x^2) //power rule for the cos(x). also you have to take the derivative of the inside of the cos(x) y' = 2*(cos(x))*(-sin(x)) + d/dx(cos(x^2)) // now take the derivative of the second part of the sum. no power rule. just take derivative of the cos() and then the derivative of the inside. y' = 2*(cos(x))*(-sin(x)) + (-sin(x^2))*(2x) // so you can factor a 2 if you want.

OpenStudy (anonymous):

Why would you have the first -sin(x)?

OpenStudy (anonymous):

Shouldn't it just be -1? I'm lost.

OpenStudy (anonymous):

you have to take the derivative of the cos(x) which is -sin(x)

OpenStudy (anonymous):

Oh right... duh. Thank you so much!

OpenStudy (anonymous):

yw